# Bounded geometry for PCF-special subvarieties

### v2

@@ Abstract.
For each integer $d\geq 2$, let $\mathrm{M}_{d}$ denote the moduli space of maps $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ of degree $d$. We study the geometric configurations of subsets of postcritically finite (or PCF) maps in $\mathrm{M}_{d}$. A complex-algebraic subvariety $Y\subset\mathrm{M}_{d}$ is said to be PCF-special if it contains a Zariski-dense set of PCF maps. Here we prove that there are only finitely many positive-dimensional irreducible PCF-special subvarieties in $\mathrm{M}_{d}$ with degree $\leq D$. In addition, there exist constants $N=N(D,d)$ and $B=B(D,d)$ so that for any complex algebraic subvariety $X\subset\mathrm{M}_{d}$ of degree $\leq D$, the Zariski closure $\overline{X\cap\mathrm{PCF}}$ has at most $N$ irreducible components, each with degree $\leq B$. We also prove generalizations of these results for points with small critical height in $\mathrm{M}_{d}(\overline{\mathbb{Q}})$.
@@ 1. Introduction
For each integer $d\geq 2$, let $\mathrm{M}_{d}$ denote the moduli space of maps $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ of degree $d$, defined over the field $\mathbb{C}$ of complex numbers. The space $\mathrm{M}_{d}$ consists of Möbius conjugacy classes of maps $f$; it is constructed as a complex orbifold in [Mi1] and as a scheme over $\operatorname{Spec}\mathbb{Z}$ in [Si1]. Let $\mathrm{PCF}\subset\mathrm{M}_{d}$ denote the subset of postcritically finite maps; i.e., the conjugacy classes of maps $f$ for which each critical point has a finite forward orbit. We say that an irreducible complex algebraic subvariety $X\subset\mathrm{M}_{d}$ is PCF-special if $\mathrm{PCF}\cap X$ is Zariski dense in $X$. Note that $\mathrm{M}_{d}$ is itself PCF-special, as is the polynomial subvariety $\mathrm{Poly}_{d}\subset\mathrm{M}_{d}$, as are all subvarieties defined by orbit relations among the critical points; see for example [BD, Proposition 2.6] and [De2, Theorem 6.2]. Conjecturally, all PCF-special subvarieties are determined by critical orbit relations (in a general sense, accounting for symmetries among the maps) [BD, Conjecture 1.10].
In this article, we study the geometry of the set $\mathrm{PCF}$ by examining how it intersects with families of varieties in $\mathrm{M}_{d}$, and we prove uniform bounds on the degrees of PCF-special subvarieties in these families. Here, and throughout this article, we measure degree with respect to a fixed choice of very ample line bundle on $\mathrm{M}_{d}$, for example providing the projective compactification of $\mathrm{M}_{d}$ defined in [Si1]. The degree bounds will depend on the choice, but the following result holds for any such choice:
@@
Consider the family of curves $\mathrm{Per}_{1}(\lambda)$ in $\mathrm{M}_{2}$, parameterized by $\lambda\in\mathbb{C}$, consisting of all quadratic maps having a fixed point of multiplier $\lambda$. This family was first introduced by Milnor in [Mi1], and they are lines in the natural coordinates that Milnor introduced on $\mathrm{M}_{2}\simeq\mathbb{C}^{2}$. It is known that $\mathrm{Per}_{1}(\lambda)$ is PCF-special if and only if $\lambda=0$ [DWY]. And although $\mathrm{PCF}$ is Zariski-dense in $\mathrm{M}_{2}$, Theorem 1.1 combined with the main result of [DWY] implies: There is a uniform bound $M$ so that
#$\#\,\mathrm{Per}_{1}(\lambda)\cap\mathrm{PCF}\;\leq\;M$
for all $\lambda\not=0$ in $\mathbb{C}$.
In fact we prove stronger versions of Theorem 1.1 and Theorem 1.2 in the spirit of the Bogomolov Conjecture [Ul, Zh1] or, more recently, of the “height gap principle” [GGK, Gao] in the setting of abelian varieties, for subvarieties of $\mathrm{M}_{d}$ defined over $\overline{\mathbb{Q}}$. The critical height of a conjugacy class $[f]\in\mathrm{M}_{d}(\overline{\mathbb{Q}})$ is defined by
[…]
for any representative $f$ defined over $\overline{\mathbb{Q}}$, where the critical points are counted with multiplicity and $\hat{h}_{f}$ is the Call-Silverman canonical height on $\mathbb{P}^{1}(\overline{\mathbb{Q}})$ [CS] [Si2]. Ingram proved that $\hat{h}_{\mathrm{crit}}$ is comparable to a Weil height on the open subset $\mathrm{M}_{d}\setminus\mathcal{L}_{d}$ where $\mathcal{L}_{d}$ is the locus of flexible Lattès maps [In]; see also [YZ, Remark 6.3.3]. The PCF points in $\mathrm{M}_{d}(\overline{\mathbb{Q}})$ are precisely those with $\hat{h}_{\mathrm{crit}}(f)=0$.
To state our results, we fix a projective embedding of $\mathrm{M}_{d}$ defined over $\mathbb{Q}$. We also fix height functions $h_{\mathrm{Ch}}$ on the associated Chow varieties $\mathrm{Ch}(\mathrm{M}_{d},r,D)$ of cycles of dimension $r$ and degree $\leq D$ in $\mathrm{M}_{d}$, for example a Philippon height [Ph]; see §2.5 for definitions. The constants $B$, $N$, and $\epsilon$ of the following theorems will depend on these choices.
@@
[…]
@@ 1.1. History and context
The study of the PCF-special varieties in $\mathrm{M}_{d}$ grew out of parallels between the theory of complex dynamical systems and the study of abelian varieties. A conjectural classification of PCF-special subvarieties in $\mathrm{M}_{d}$ first appears in [Si2, Chapter 6] where it was given the name of “Dynamical André-Oort” (or “DAO” as it is known in [JX]), because the geometry of the PCF locus in $\mathrm{M}_{d}$ shares features with the locus of CM points in a moduli space of abelian varieties; see [Za] for background on several Unlikely Intersection problems. The first cases of DAO were proved by Baker-DeMarco [BD] and Ghioca-Hsia-Tucker [GHT], followed by a series of results mostly focused on polynomial dynamics where the relations and symmetries are understood, such as [GY, FG2], culminating in a complete classification of PCF-special curves in moduli spaces of polynomials by Favre-Gauthier [FG3] and a characterization of PCF-special curves in $\mathrm{M}_{d}$ by Ji-Xie [JX]. We note, however, that there are currently no known classification results for PCF-special subvarieties of dimensions $>1$, even when restricting to spaces of polynomials. Nevertheless, Theorem 1.1 can be viewed as a “Uniform DAO” that holds for PCF-special subvarieties of any dimension.
@@
A broader conjecture in [DM3] contains the classification of PCF-special subvarieties (as formulated in [BD, Conjecture 1.10]) and the uniform bounds on their geometry (our Theorems 1.1 and 1.2) and many results about geometry of torsion points in abelian varieties as special cases. It also contains the (partial) uniformity results obtained in [DM1] and [GTV].
@@ 1.2. Proof strategy
For any positive integer $m$, we let $\mathcal{X}_{V}^{m}$ denote the $m$-th fiber power of $\mathcal{X}$ over $V$. We denote elements of $\mathcal{X}_{V}^{m}$ by $(\lambda,f_{1},\ldots,f_{m})$, with $f_{i}\in X_{\lambda}$ for each $i$. Our main theorems are derived from the following result.
@@
The contradiction will come from an analysis of the two measures. More precisely, using a slicing argument as in [MaS] (which was revisited in [DM1]), it suffices to analyze the family of bifurcation measures they induce on the fibers $X_{\lambda}$. Here we appeal to the maximal variation of the family and the dynamical theory of $J$-stability. We find a parameter $\lambda_{0}\in V$ and conjugacy class $[f_{0}]\in X_{\lambda_{0}}$ so that one collection of $r$ critical points is pre-repelling for $f_{0}$, while one of the critical points in the other $r$-tuple is in the basin of a parabolic point. We then zoom into the parameter space at $f_{0}$, where the first $r$-tuple of preperiodic critical points yields an asymptotic similarity between the bifurcation measure and the dynamical space, exactly as in the proof of [BE, Main Theorem]. But zooming in at this rate yields 0 for the bifurcation measure defined with the second $r$-tuple of critical points, because one of the critical points has infinite orbit and is attracted to a parabolic cycle. Details are provided in Section 4. The proofs of Theorems 1.1, 1.2, 1.5, and 1.6 are completed in Section 5.
@@ 1.3. Questions
on $\Lambda$. Can we conclude that there exists $i\in\{1,\ldots,r-1\}$ so that the pair $(c_{i},c_{r+1})$ is dynamically related along $\Lambda$? (We remark here that one can define a dynamically related tuple $(c_{1},\ldots,c_{r+1})$ similarly; see [De2, §6.2]. But it follows by [MeSc, Proposition 2.21], that dynamical relations always come in pairs.)
@@ 1.4. Acknowledgements
We would like to thank Dori Bejleri, Xavier Buff, Ziyang Gao, Thomas Gauthier, Zhuchao Ji, Lars Kühne, Harry Schmidt, Gabriel Vigny, and Junyi Xie for helpful discussions. Our research was supported in part by the National Science Foundation and by the Natural Sciences and Engineering Research Council of Canada. In addition, Laura acknowledges the Radcliffe Institute for Advanced Study at Harvard; Myrto acknowledges the University of Toronto and Fields Insitute; and Hexi acknowledges Zhejiang University for their support of this collaboration.
@@ 2.1. The moduli space $\mathrm{M}_{d}$ and a critically-marked parameter space
For each degree $d\geq 2$, the moduli space $\mathrm{M}_{d}$ is the space of conformal conjugacy classes of maps $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ defined over $\mathbb{C}$. It is an affine algebraic variety of dimension $22d-$ 2. Silverman showed that it exists as a geometric quotient over $\operatorname{Spec}\mathbb{Z}$, and he introduced a projective compactification $\overline{M}_{d}$ via the tools of Geometric Invariant Theory which can be defined over $\mathbb{Q}$ [Si1]. We note that $\mathrm{M}_{2}\simeq\mathbb{C}^{2}$ [Mi1] while $\mathrm{M}_{d}$ is singular in all degrees $d>2$ [We, MSW]. Because of the existence of maps with automorphisms, this $\mathrm{M}_{d}$ is not a fine moduli space, but we can pass to a finite branched cover $\mathcal{M}_{d}^{cm}\to\mathrm{M}_{d}$ to build a family
@@ 2.2. PCF maps and Thurston rigidity
A map $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ is postcritically finite if each of its critical points has a finite forward orbit. An important class of examples of such maps are the Lattès maps, which are quotients of a self-map of a complex torus; the flexible Lattès maps are those that are part of a positive-dimensional family in $\mathrm{M}_{d}$; see [Mi2] for more information on Lattès maps. We will denote the flexible Lattès locus by $\mathcal{L}_{d}$ in $\mathrm{M}_{d}$; it exists only in square degrees. From Thurston’s rigidity theorem [DH], we may deduce that any postcritically finite map which is not flexible Lattès is Möbius-conjugate to a map $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ defined over $\overline{\mathbb{Q}}$.
@@ 2.3. The bifurcation current and its wedge powers
$T_{i}=\pi_{*}\left(\hat{T}_{f}\wedge[\Gamma_{c_{i}}]\right),$
where $\pi:\mathcal{M}_{d}^{cm}\times\mathbb{P}^{1}\to\mathcal{M}_{d}^{cm}$ is the projection and $[\Gamma_{c_{i}}]$ is the current of integration along the graph of $c_{i}$. Then $T_{i}$ is a positive (1,1)-current with continuous potentials. The sum
$T_{\mathrm{bif}}=\sum_{i=1}^{2d-2}T_{i}$

### v3

@@ Abstract.
For each integer $d\geq 2$, let $\mathrm{M}_{d}$ denote the moduli space of maps $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ of degree $d$. We study the geometric configurations of subsets of postcritically finite (or PCF) maps in $\mathrm{M}_{d}$. A complex-algebraic subvariety $Y\subset\mathrm{M}_{d}$ is said to be PCF-special if it contains a Zariski-dense set of PCF maps. Here we prove that there are only finitely many positive-dimensional irreducible PCF-special subvarieties in $\mathrm{M}_{d}$ with degree $\leq D$. In addition, there exist constants $N=N(D,d)$ and $B=B(D,d)$ so that for any complex algebraic subvariety $X\subset\mathrm{M}_{d}$ of degree $\leq D$, the Zariski closure $\overline{X\cap\mathrm{PCF}_{d}}$ has at most $N$ irreducible components, each with degree $\leq B$. We also prove generalizations of these results for points with small critical height in $\mathrm{M}_{d}(\overline{\mathbb{Q}})$.
@@ 1. Introduction
For each integer $d\geq 2$, let $\mathrm{M}_{d}$ denote the moduli space of maps $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ of degree $d$, defined over the field $\mathbb{C}$ of complex numbers. The space $\mathrm{M}_{d}$ consists of Möbius conjugacy classes of maps $f$; it is constructed as a complex orbifold in [Mi1] and as a scheme over $\operatorname{Spec}\mathbb{Z}$ in [Si1]. Let $\mathrm{PCF}_{d}\subset\mathrm{M}_{d}$ denote the subset of postcritically finite maps; i.e., the conjugacy classes of maps $f$ for which each critical point has a finite forward orbit. We say that an irreducible complex algebraic subvariety $X\subset\mathrm{M}_{d}$ is PCF-special if $\mathrm{PCF}_{d}\cap X$ is Zariski dense in $X$. Note that $\mathrm{M}_{d}$ is itself PCF-special, as is the polynomial subvariety $\mathrm{MPoly}_{d}\subset\mathrm{M}_{d}$, as are all subvarieties defined by orbit relations among the critical points; see for example [BD, Proposition 2.6] and [De2, Theorem 6.2]. Conjecturally, all PCF-special subvarieties are determined by critical orbit relations (in a general sense, accounting for symmetries among the maps) [BD, Conjecture 1.10].
In this article, we study the geometry of the set $\mathrm{PCF}_{d}$ by examining how it intersects with families of varieties in $\mathrm{M}_{d}$, and we prove uniform bounds on the degrees of PCF-special subvarieties in these families. Here, and throughout this article, we measure degree with respect to a fixed choice of very ample line bundle on $\mathrm{M}_{d}$, for example providing the projective compactification of $\mathrm{M}_{d}$ defined in [Si1]. The degree bounds will depend on the choice, but the following result holds for any such choice:
@@
Consider the family of curves $\mathrm{Per}_{1}(\lambda)$ in $\mathrm{M}_{2}$, parameterized by $\lambda\in\mathbb{C}$, consisting of all quadratic maps having a fixed point of multiplier $\lambda$. This family was first introduced by Milnor in [Mi1], and they are lines in the natural coordinates that Milnor introduced on $\mathrm{M}_{2}\simeq\mathbb{C}^{2}$. It is known that $\mathrm{Per}_{1}(\lambda)$ is PCF-special if and only if $\lambda=0$ [DWY]. And although $\mathrm{PCF}_{d}$ is Zariski-dense in $\mathrm{M}_{2}$, Theorem 1.1 combined with the main result of [DWY] implies: There is a uniform bound $N$ so that
$\mathrm{Per}_{1}(\lambda)\cap\mathrm{PCF}_{d}\;\leq\;N$
for all $\lambda\not=0$ in $\mathbb{C}$. Theorem 1.1 also implies the corresponding result for the family of curves $\mathrm{Per}_{n}(\lambda)$ in the moduli space $\mathrm{MPoly}_{3}$ of cubic polynomials, for any fixed $n\geq 1$, in view of [BD, Theorem 1.1] [FG2, Theorem B].
In fact we prove stronger versions of Theorem 1.1 and Theorem 1.2 in the spirit of the Bogomolov Conjecture [Ul, Zh1] or, more recently, of the “height gap principle” [GGK, Gao] in the setting of abelian varieties, for subvarieties of $\mathrm{M}_{d}$ defined over $\overline{\mathbb{Q}}$. The critical height of a conjugacy class $f]\in\mathrm{M}_{d}(\overline{\mathbb{Q}})$ is defined by
$\hat{h}_{\mathrm{crit}}(f)=\sum_{\{c~:~f^{\prime}(c)=0\}}\hat{h}_{f}(c)\;\geq\;0,$
for any representative $f$ defined over $\overline{\mathbb{Q}}$, where the critical points are counted with multiplicity and $\hat{h}_{f}$ is the Call-Silverman canonical height on $\mathbb{P}^{1}(\overline{\mathbb{Q}})$ [CS] [Si2]. Ingram proved that $\hat{h}_{\mathrm{crit}}$ is comparable to a Weil height on the open subset $\mathrm{M}_{d}\setminus\mathcal{L}_{d}$ where $\mathcal{L}_{d}$ is the locus of flexible Lattès maps [In]. A more explicit version of Ingram’s result was later established in [GOV, Theorem C]. A different approach is given by the Yuan–Zhang theory of adelic line bundles; see [YZ, Remark 6.3.3]. The PCF points in $\mathrm{M}_{d}(\overline{\mathbb{Q}})$ are precisely those with $\hat{h}_{\mathrm{crit}}(f)=0$.
To state our results, we fix a projective embedding of $\mathrm{M}_{d}$ defined over $\mathbb{Q}$. We also fix height functions $h_{\mathrm{Ch}}$ on the associated Chow varieties $\mathrm{Ch}(\mathrm{M}_{d},r,D)$ of cycles of dimension $r$ and degree $\leq D$ in $\mathrm{M}_{d}$, for example a Philippon height [Ph]; see §2.5 for definitions. The constants $B$, NN, and $\epsilon$ of the following theorems will depend on these choices.
@@
$\mathrm{ess.min.}(\hat{h}_{\mathrm{crit}}|_{Y})\;:=\;\sup_{Y^{\prime}\subset Y}\inf_{y\in(Y\setminus Y^{\prime})(\overline{\mathbb{Q}})}\hat{h}_{\mathrm{crit}}(y)\;\geq\;0,$
@@ 1.1. History and context
The study of the PCF-special varieties in $\mathrm{M}_{d}$ grew out of parallels between the theory of complex dynamical systems and the study of abelian varieties. A conjectural classification of PCF-special subvarieties in $\mathrm{M}_{d}$ first appears in [Si2, Chapter 6] where it was given the name of “Dynamical André-Oort” (or “DAO” as it is known in [JX]), because the geometry of the PCF locus in $\mathrm{M}_{d}$ shares features with the locus of CM points in a moduli space of abelian varieties; see [Za] for background on several Unlikely Intersection problems. The first cases of DAO were proved by Baker-DeMarco [BD] and Ghioca-Hsia-Tucker [GHT], followed by a series of results mostly focused on polynomial dynamics where the relations and symmetries are understood, such as [GY, FG2], culminating in a complete classification of PCF-special curves in moduli spaces of polynomials by Favre-Gauthier [FG3] and a characterization of PCF-special curves in $\mathrm{M}_{d}$ by Ji-Xie [JX]. We note, however, that there are currently no known classification results for PCF-special subvarieties of dimensions >1>1, even when restricting to spaces of polynomials. Nevertheless, Theorem 1.1 can be viewed as a “Uniform DAO” that holds for PCF-special subvarieties of any dimension.
@@
A far-reaching conjecture in [DM3] implies the classification of PCF-special subvarieties (as formulated in [BD, Conjecture 1.10]) and the uniform bounds on their geometry (our Theorems 1.1 and 1.2) and many results about geometry of torsion points in abelian varieties as special cases. It also contains the (partial) uniformity results obtained in [DM1] and [GTV].
@@ 1.2. Proof strategy
For any positive integer $m$, we let $\mathcal{X}_{V}^{m}$ denote the $m$-th fiber power of $\mathcal{X}$ over $V$. We denote elements of $\mathcal{X}_{V}^{m}$ by $(\lambda,f_{1},\ldots,f_{m})$, with $f_{i}\in X_{\lambda}$ for each $i$. Our main theorems are derived from the following result, which is inspired from the Relative Bogomolov Conjecture [DGH2] and by results in [] and [GGK].
@@
The contradiction will come from an analysis of the two measures. More precisely, using a slicing argument as in [MaS] (which was revisited in [DM1]), it suffices to analyze the family of bifurcation measures they induce on the fibers $X_{\lambda}$. Here we appeal to the maximal variation of the family and the dynamical theory of $J$-stability. We find a parameter $\lambda_{0}\in V$ and conjugacy class $f_{0}]\in X_{\lambda_{0}}$ so that one collection of $r$ critical points is pre-repelling for $f_{0}$, while one of the critical points in the other $r$-tuple is in the basin of a parabolic point. We then zoom into the parameter space at $f_{0}$, where the first $r$-tuple of preperiodic critical points yields an asymptotic similarity between the bifurcation measure and the dynamical space, exactly as in the proof of [BE, Main Theorem]. But zooming in at this rate yields 0 for the bifurcation measure defined with the second $r$-tuple of critical points, because one of the critical points has infinite orbit and is attracted to a parabolic cycle. Details are provided in Section 4. The proofs of Theorems 1.1, 1.2, 1.5, and 1.6 are completed in Section 5.
@@ 1.3. Questions
on $\Lambda$. Can we conclude that there exists $i\in\{1,\ldots,r-1\}$ so that the pair $(c_{i},c_{r+1})$ is dynamically related along $\Lambda$? (We remark here that one can define a dynamical relation for a tuple $(c_{1},\ldots,c_{r+1})$ as for a pair; see [De2, §6.2]. But it follows by [MeSc, Proposition 2.21] that dynamical relations always come in pairs.) We know of no cases of this question that have been answered.
Remark 1.11.
Versions of Theorems 1.1 and 1.2 are expected to hold in more general settings, for example when marking points on $\mathbb{P}^{1}$ are not necessarily critical, or upon working with preperiodic points in families of maps on $\mathbb{P}^{N}$. See [DM3]. Analogues of Questions 1.9 and 1.10 can also be formulated, but we keep our focus in this article on the postcritically-finite maps on $\mathbb{P}^{1}$ and the bifurcation currents associated to marked critical points.
@@ 1.4. Acknowledgements
We would like to thank Dori Bejleri, Xavier Buff, Ziyang Gao, Thomas Gauthier, Zhuchao Ji, Lars Kühne, Harry Schmidt, Gabriel Vigny, and Junyi Xie for helpful discussions. We also thank the anonymous referees for their comments and suggestions. Our research was supported in part by the National Science Foundation and by the Natural Sciences and Engineering Research Council of Canada. In addition, Laura acknowledges the Radcliffe Institute for Advanced Study at Harvard; Myrto acknowledges the University of Toronto and Fields Insitute; and Hexi was partially supported by National Key R&D Program of China (20212021YFA10032021003202) and NSFC ($12131016$, $12331004$).
@@ 2.1. The moduli space $\mathrm{M}_{d}$ and a critically-marked parameter space
For each degree $d\geq 2$, the moduli space $\mathrm{M}_{d}$ is the space of conformal conjugacy classes of maps $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ defined over $\mathbb{C}$. It is an affine algebraic variety of dimension 2d22d-2. Silverman showed that it exists as a geometric quotient over $\operatorname{Spec}\mathbb{Z}$, and he introduced a projective compactification $\overline{M}_{d}$ via the tools of Geometric Invariant Theory which can be defined over $\mathbb{Q}$ [Si1]. We note that $\mathrm{M}_{2}\simeq\mathbb{C}^{2}$ [Mi1] while $\mathrm{M}_{d}$ is singular in all degrees $d>2$ [We, MSW]. Because of the existence of maps with automorphisms, this $\mathrm{M}_{d}$ is not a fine moduli space, but we can pass to a finite branched cover $\mathcal{M}_{d}^{cm}\to\mathrm{M}_{d}$ to build a family
@@ 2.2. PCF maps and Thurston rigidity
A map $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ is postcritically finite if each of its critical points has a finite forward orbit. An important class of examples of such maps are the Lattès maps, which are quotients of a self-map of a complex torus; the flexible Lattès maps are those that are part of a positive-dimensional family in $\mathrm{M}_{d}$; see [Mi2] for more information on Lattès maps. We will denote the flexible Lattès locus by $\mathcal{L}_{d}$ in $\mathrm{M}_{d}$; it exists only in square degrees. From Thurston’s rigidity theorem [Douady:Hubbard:Thurston], we may deduce that any postcritically finite map which is not flexible Lattès is Möbius-conjugate to a map $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ defined over $\overline{\mathbb{Q}}$.
@@ 2.3. The bifurcation current and its wedge powers
(2.2) $T_{i}=\pi_{*}\left(\hat{T}_{f}\wedge[\Gamma_{c_{i}}]\right),$
where $\pi:\mathcal{M}_{d}^{cm}\times\mathbb{P}^{1}\to\mathcal{M}_{d}^{cm}$ is the projection and $\Gamma_{c_{i}}]$ is the current of integration along the graph of $c_{i}$. Then $T_{i}$ is a positive (1,1)-current with continuous potentials. The sum
(2.3) $T_{\mathrm{bif}}=\sum_{i=1}^{2d-2}T_{i}$